DIFFERENTIAL GEOMETRY/PDE SEMINAR

Wednesday, May 22, 2002 Thomson 134 3:45 pm

Poincaré Metrics and Conformal Geometry

Robin Graham (University of Washington)

I will describe some of the recently uncovered connections between asymptotic behavior at infinity on a conformally compact asymptotically Einstein manifold and conformally invariant objects on its boundary at infinity.

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.